OpenStudy (anonymous):

Verify that f and g are inverse functions. f(x)= 2x + 5 g(x)= x -5/ 2

OpenStudy (anonymous):

plug y into one of the equations, if it comes out identical to the other then theyre reverse functions

OpenStudy (anonymous):

basically just choose any one of the two equations and switch y and x around

OpenStudy (radar):

It is said, "I don't see a "y"" Maybe f(x) is y. If so y=2x+5 Now switch the y and x x=2y+5 2y=x-5 y=x/2 - 5/2 I don't know, but that don't look the same as g(x)

OpenStudy (radar):

Maybe, I chose the wrong equation, lets look at the other one g(x) = x -5/2 y=x-5/2 x=y-5/2 y=x+5/2 doesn't look like 2x+5. Don't think they are inverse functions.